OpenStudy (anonymous):

√48a^5b divided by √12ab

OpenStudy (anonymous):

Are you sure that's right?

OpenStudy (anonymous):

√48a^5b 4√ 3 a^4 ------- = ------- = 2a^4 √12ab 2√3

OpenStudy (anonymous):

Yeah that looks more like it!

OpenStudy (anonymous):

oh - i could have done it easier - because √48 / √12 = √4 = 2

OpenStudy (anonymous):

Could you answer some more for me please?

OpenStudy (anonymous):

\[\sqrt{\frac{48 a^5 b}{12a b}}=2 a^2 \]

OpenStudy (anonymous):

yeah - but please post them separately - if theres a few post them together

OpenStudy (anonymous):

oh rob may be right - does the √ reger to just the number or all of the expression - i assumed it was just the number

OpenStudy (anonymous):

*refer

OpenStudy (anonymous):

@jimmyrep Yes, it's another ambiguous radical symbol. Assumed that it covered the entire expression length.

OpenStudy (anonymous):

yea - i usually ask the person first - ah well!! we both get kudos now!!

OpenStudy (anonymous):

@jimmyrep I usually don't answer problems like this one where there is a question regarding the presentation.